Light travelling from water into crowd glass, strikes the crown glass at an angle of 41.7 degrees. What is the angle of refraction?
1
Expert's answer
2020-12-03T10:59:31-0500
"\frac{\sin{41.7\degree}}{n_1}=\frac{\sin{a}}{n_2}\\a=\arcsin{\frac{1.52\sin{41.7\degree}}{1.33}}=49.5\degree"